# FORGE: Force-Guided Exploration for Robust Contact-Rich Manipulation under Uncertainty

### v1

@@ Abstract
We present FORGE, a method that enables sim-to-real transfer of contact-rich manipulation policies in the presence of significant pose uncertainty. FORGE combines a force threshold mechanism with a dynamics randomization scheme during policy learning in simulation, to enable the robust transfer of the learned policies to the real robot. At deployment, FORGE policies, conditioned on a maximum allowable force, adaptively perform contact-rich tasks while respecting the specified force threshold, regardless of the controller gains. Additionally, FORGE autonomously predicts a termination action once the task has succeeded. We demonstrate that FORGE can be used to learn a variety of robust contact-rich policies, enabling multi-stage assembly of a planetary gear system, which requires success across three assembly tasks: nut-threading, insertion, and gear meshing. Project website: https://noseworm.github.io/forge/
@@ I Introduction
We are interested in developing sim-to-real techniques for learning assembly primitives (e.g., low-clearance insertion or nut-threading). Over the past decade, work in simulation and sim-to-real techniques has led to advances in challenging areas such as dexterous manipulation and legged locomotion [1, 2, 3, 4]. However, similar results have only recently been achieved for robotic assembly, which requires efficient and accurate simulation of both the robot and the detailed, low-clearance parts [5, 6, 7, 8, 9, 10].
Even with these advances, successfully deploying sim-to-real policies for assembly tasks remains challenging. Previous approaches typically only consider small amounts of perceptual noise relative to part size. This assumption aligns with industrial-style robot workcells where uncertainty is typically engineered away. Strategies to deal with uncertainty include careful mechanical design of fixtures and adapters, extensive calibration processes, and the use of high-precision sensing. As a result, the time and money required to set up a workcell with different parts, poses, and tasks can be prohibitive for small-scale enterprises, or in less-structured environments. We aim to develop control methods that are robust to higher levels of pose estimation error, which is unavoidable in less structured environments.
When there is pose uncertainty, behaviours that search for and rely on contact can be used to ensure success [11, 12]. However, the required contact between the parts can lead to undesirable outcomes if the force is too high. Parts can slip or become damaged, making the task difficult or impossible to complete. Previous heuristic approaches, such as spiral search [11, 13], can limit the applied force but are task-specific and can be inefficient. RL offers a general paradigm for developing more flexible search behaviours. However, the sim-to-real gap makes it difficult to transfer policies learned in simulation to the real world. In particular, it is challenging to simulate a similar contact distribution to what the real robot would experience. Even if the simulator has an accurate robot model (itself a time-consuming calibration procedure), it is difficult to know a priori the material and inertial properties of the parts the robot will interact with.
@@
Figure 1: FORGE uses force feedback to learn search behaviours for contact-rich tasks with pose estimation uncertainty. It combines dynamics randomization, a force threshold, and early termination for robust sim-to-real transfer. The resulting policies are safe and efficient (bottom) compared to aggressive baseline policies that cause parts to slip (top).
@@
In this work, we propose FORGE: a framework for developing sim-to-real policies that safely and efficiently perform assembly tasks in the presence of significant pose uncertainty. FORGE trains policies in simulation that are robust to a wide range of contact interactions. Additionally, policies are trained without precise knowledge of part poses, leading to emergent search behaviours.
@@
FORGE has two complementary components to ensure policies are robust to contact. First, we propose to condition policies on a force threshold that should not be exceeded during task execution. Second, policies are trained to maintain this threshold under a wide range of dynamics randomizations (we randomize robot, controller, and part properties). The dynamics randomization ensures that the search behaviour is robust to simulator mis-specification and has the added benefit that the resulting policies can be deployed without extensive controller tuning.
As assembly policies become more performant, we emphasize the importance of reporting metrics beyond success rate, such as mean force or time to success. Standard practice in sim-to-real assembly is to execute policies for a fixed duration [7]. However, due to task variation, this will often lead to premature termination or leave a robot “waiting” after the task is finished. Instead, the policy can determine when to terminate. This is itself a difficult task that can benefit from contact (e.g., a successfully inserted peg cannot move laterally). FORGE proposes a method for early termination that expands the action space so that the policy learns to predict task success. We show that early termination, also trained in simulation, robustly transfers to the real world.
@@
    A method to specify maximum allowable contact-force during policy execution. This results in policies that exhibit safe search behaviour even with significant levels of pose estimation error (up to $5mm$).
@@
    A dynamics randomization scheme that enables robust sim-to-real transfer, and minimizes the need to tune controller gains.
@@
    A method for early termination prediction that allows efficient policy execution.
@@
Results are shown over $>500$ real-world trials and multiple tasks. We plan to release the code with the paper.
@@ II-A Assembly Tasks
Figure 2: FORGE is evaluated on three tasks proposed in the Factory work [5]: Peg Insertion, Gear Meshing, and Nut Threading. Each task is trained solely in simulation (top) and transferred directly to the real robot (bottom).
Each task involves mating two parts: one held in the gripper and another fixed to the workspace. We consider all three tasks from Factory [5] and demonstrate the first sim-to-real transfer for threading a small M16 nut (see Fig. 2).
Peg Insertion: A small round peg with $8mm$ diameter needs to be inserted into a corresponding socket with $0.5mm$ diametrical clearance. There is position uncertainty such that a successful search behaviour requires lateral exploration.
Gear Meshing: Gears need to be inserted onto pegs with $0.5mm$ clearance. Other gears are present and the teeth of adjacent gears must be aligned for successful meshing. In addition to lateral exploration, rotational exploration may be necessary to mesh the teeth.
@@ II-B POMDP Formulation
We formulate our problem as a Partially Observable Markov Decision Process (POMDP) [15, 16] to reflect the partial observability of most contact-rich manipulation setups. A POMDP is defined by the tuple: $(\mathcal{S},\mathcal{A},\Omega,T,O,R,\gamma$). The goal is to learn a parameterized policy, $\pi_{\theta}(a_{t}|o_{1},\ldots,o_{t})$, that maximizes the expected return:
$J(\pi_{\theta})=\mathbb{E}_{\tau\sim p(\tau,|\pi_{\theta},\Psi)}[\Sigma_{t=0}^{\infty}\gamma^{t}r_{t}]$ (1)
@@
[…]
Observations ($\Omega$): It is difficult to accurately estimate the full state for contact-rich manipulation. Instead, all our policies use the following observations:
@@
We do not include pose or velocity of the held part because it can move in the gripper and be difficult to track without tactile sensing. Likewise, we do not observe $\Psi$, but include the previous action, $a_{t-1}$, to help infer unknown dynamics.
@@
Transition Function ($T$): $T$ determines the next state and is parameterized by the dynamics parameters, $\Psi$: $T_{\Psi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$. $T_{\Psi}$ is often specified using a simulator (in our case IsaacGym [18]) with the corresponding set of simulation parameters, $\Psi^{sim}$. The sim-to-real gap comes from the mismatch between $\Psi^{sim}$ and $\Psi^{real}$.
Observation Function ($O$): The observation function generates noisy observations from state: $O:\mathcal{S}\times\mathcal{A}\rightarrow\Omega$. The position of the fixed part is assumed to have up to $5mm$ error. Independent Gaussian noise is added to each of the other observations at every timestep (except velocity, where positional noise is propagated through finite differencing). Noise values can be found in Appendix -A.
Reward ($R$): The reward function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, uses a keypoint formulation as its main component: $R_{kp}(p^{fixed},p^{held}_{t})$. The target keypoints, $k^{targ}$, represent the desired position of the held part, while $k^{held}_{t}$ represent its current position. We modify the keypoint reward from previous work [5, 19] to account for small, threaded geometries (see Appendix -B for more details). We also add two discrete bonus rewards that are given when important phases of the tasks are reached: once the held part is centered on top of the fixed part and once the task is successful:
@@
The relative $z$-position of bottom of the held part to the top of the fixed part is used to check each condition (see Appendix -B). We found the bonuses led to more robust learning when there is significant pose uncertainty.
@@ III FORGE: Robust Search under Uncertainty
FORGE uses on-policy RL to learn exploratory behaviours in simulation. A force threshold (Section III-A) and dynamics randomization (Section III-B) are introduced to achieve robust search behaviours. FORGE also introduces an early termination procedure (Section III-C) for efficient execution.
@@ III-A Force Threshold
During policy execution, excessive force can cause parts to slip or become damaged (e.g., electronic components with fragile pins). Although it may be possible to recover from small amounts of slip with the right sensors (e.g., wrist camera or tactile), we prefer to avoid these scenarios.
To develop safe policies, we propose to condition the policy on a force threshold, $F_{th}$: $\pi(a|o,F_{th})$. During training, the policy is penalized if the contact force, $F^{ee}_{t}$, experienced by the arm exceeds the threshold. Concretely, we add an additional term to the reward function:
$R_{contact\_pen}(F^{ee}_{t})=-\beta*\max(0,||F^{ee}_{t}||-F_{th}).\vspace{-0.5em}$ (3)
In simulation, the true contact force can be measured. Note that this penalty can be used during training whether the policy has access to the force observation or not.
@@ III-B Dynamics Randomization
To successfully deploy policies trained in simulation, it is important that the trajectory distribution experienced during training is similar to what it would be when deployed: $p(\tau^{real}|\pi_{\theta},\Psi^{real})\approx p(\tau^{sim}|\pi_{\theta},\Psi^{sim})$. The difference between these distributions is usually referred to as the sim-to-real gap. To gain insight into why minimizing the gap is important, specifically for contact-rich tasks, we can look more deeply into how trajectories are sampled:
$\begin{split}\tau\sim p(\tau|\pi,\Psi)=&p(s_{0})\prod\limits_{t=1}^{T}\Big{[}\pi(a_{t}|o_{1:t})\\ &p(o_{t}|s_{t},a_{t-1},\Psi)p(s_{t}|s_{t-1},a_{t-1},\Psi)\Big{]}.\end{split}$ (4)
From this equation, we see the dynamics parameters can impact both the next-state and observation distributions. For the same action, different dynamics parameters can lead to parts being in different locations. Further, similar actions may lead to different observed contact forces.
The sim-to-real gap is usually handled by (1) system identification (Sys-ID) [20] or (2) dynamics randomization (DR) [21, 10]. The goal of Sys-ID is to tune $\Psi^{sim}$ to be close to $\Psi^{real}$. This itself is a complicated tuning procedure that may need to be redone for every new set of parts. Instead, we follow the DR approach which learns policies that are robust to a wide range of dynamics parameters. Concretely, we optimize a version of Eq. 1 where:
@@
The integral is approximated with Monte Carlo samples from a randomization distribution (see Appendix -A for values). The next subsections describe the variables that are randomized (at the beginning of each episode).
Controller Randomization: The controller has a large impact on what force will be experienced. This work uses impedance-control where applied forces are computed as:
@@
First, the policy outputs a relative-pose, $a_{t}$, which is applied to the fixed part’s pose to get an absolute target pose, $p^{targ}_{t}$. This pose is clipped by an action scale, $\lambda$, to ensure that the target is not too far from the EE’s current pose. As in previous work, we use critically damped gains to ensure stable controllers: $k_{d}=2\sqrt{k_{p}}$ [10, 22, 23]. The controller thus depends on two parameters which govern how much force can be commanded: $\lambda\times k_{p}$. We randomize both quantities so that the range of maximum commandable forces is in $[6.4,20.0]N$. Note that the control parameters are not included in the observations, so the policy must adjust its behavior based on force measurements. This reduces the policy’s dependence on a particular controller implementation.
Controller tuning [7] or optimization [23] is a costly and often complex procedure. Randomization has the additional benefit that the policy is robust to a range of control parameters, greatly simplifying deployment.
Part Friction Randomization: As parts slide against each other, the material friction determines how much lateral force the sensor will experience. To ensure policies can work across a range of materials, we randomize part friction.
Robot Dynamics Randomization: Due to phenomena such as joint friction, the applied force may be smaller than the commanded force. We implement a simple way to account for this: inducing a randomized dead-zone in simulation. Each episode, a dead-zone is selected for each dimension, $F^{DZ}_{i}$, where commanded forces below this value are clamped to zero: $|F^{applied}_{i}|=\max(0,|F^{targ}_{i}|-F^{DZ}_{i})$. This enables the policy to increase its target which can help apply more force when needed or reduce steady-state error.
@@
III-C Early Termination
Ideally, we want the policy to terminate as soon as the task has succeeded and no sooner. Although success is clearly defined in simulation where we have access to the positions of each part, it is much more difficult to reliably predict success in the real world [24]. Consider the nut-threading task, where the distance between a successfully threaded nut and a loose nut is a fraction of a millimeter.
We propose to train a success predictor which can robustly transfer from sim-to-real and be used to make early termination decisions. Concretely, we share the weights of the policy network with the success predictor by expanding the action space of the policy to include an early termination action: $a^{ET}_{t}\in[0,1]$. To train the policy to output the correct action, we include an early termination reward, $R^{ET}_{t}$, which penalizes incorrect success predictions:

### v2

@@ Abstract
We present FORGE, a method for sim-to-real transfer of force-aware manipulation policies in the presence of significant pose uncertainty. During simulation-based policy learning, FORGE combines a force threshold mechanism with a dynamics randomization scheme to enable robust transfer of the learned policies to the real robot. At deployment, FORGE policies, conditioned on a maximum allowable force, adaptively perform contact-rich tasks while avoiding aggressive and unsafe behaviour, regardless of the controller gains. Additionally, FORGE policies predict task success, enabling efficient termination and autonomous tuning of the force threshold. We show that FORGE can be used to learn a variety of robust contact-rich policies, including the forceful insertion of snap-fit connectors. We further demonstrate the multistage assembly of a planetary gear system, which requires success across three assembly tasks: nut threading, insertion, and gear meshing. Project website: https://noseworm.github.io/forge/
@@ I Introduction
We are interested in developing sim-to-real techniques for learning assembly primitives (e.g., low-clearance insertion or nut-threading). Over the past decade, sim-to-real techniques have led to advances in dexterous manipulation and legged locomotion [1, 2, 3, 4]. However, similar results have only recently been achieved for robotic assembly, which requires efficient and accurate simulation of the detailed, low-clearance parts [5, 6, 7, 8, 9, 10]. Even with these advances, successful sim-to-real deployment remains challenging for contact-rich tasks.
Naively, policies can be too aggressive, leading to catastrophic part slip or damage that makes the task difficult or impossible to complete (see Figure 1). This is particularly pronounced when there is pose uncertainty and search behaviours that rely on contact are necessary [11, 12]. The required contact between parts can lead to undesirable outcomes if the forces are too high. Heuristic approaches, such as spiral search [11, 13], can limit the applied force but these approaches are task-specific and can be inefficient.
Reinforcement learning offers a general paradigm for developing more flexible search behaviours. However, previous works typically rely on additional procedures to ensure policies are deployed safely with desirable force profiles. For example, policies trained using the IndustReal framework [7] do not observe or adapt to contact forces. Instead, as our experiments show, forceful behaviour is determined by careful controller design and gain tuning. Other works provide methods to optimize or adapt gains online [14, 15]. Importantly, the desired force profile of a policy depends on the task at hand. For example, threading a nut may fail if the applied force is too high, while a snap-fit connector might require large forces to ensure proper insertion. Therefore, it is important to have simple and efficient methods to tune the policy’s force profile.
@@
Figure 1: FORGE uses force feedback to learn search behaviours for contact-rich tasks with position estimation uncertainty. It combines dynamics randomization, a force threshold, and success prediction for robust sim-to-real transfer. The resulting policies are safe and efficient (bottom) compared to aggressive baseline policies that cause parts to slip (top).
@@
In this work, we propose FORGE: a framework for developing force-aware sim-to-real policies for assembly tasks. FORGE policies, trained solely in simulation, use external force observations to achieve efficient and gentle behaviour. Additionally, policies are trained without precise knowledge of part poses, leading to emergent search behaviours that are robust to significant levels of pose uncertainty.
@@
FORGE has two complementary components to ensure policies are robust to contact. First, we propose to condition policies on a force threshold that should not be exceeded during task execution. Second, policies are trained to maintain this threshold under a wide range of dynamics randomizations (we randomize robot, controller, and part properties). Together, these components result in policies that can modulate their actions to achieve a force profile that respects the interpretable scalar force threshold. By randomizing this threshold during training, we are able to tune it at deployment time without retraining the policy.
For tasks with high enough clearance, a small force threshold is sufficient and no tuning is necessary. However, tasks that require significant force to succeed (e.g., snap-fit connectors) may fail if the threshold is set too low. When the required force is not known a priori, we present an automatic tuning procedure which leverages a notion of success prediction. Based on the outcome of a policy execution, the threshold can be iteratively adjusted for future trials. To automate this tuning, FORGE policies are trained to predict whether an episode succeeded or failed. We validate this procedure by showing successful sim-to-real transfer on a snap-fit connector requiring $15N$ for insertion.
Furthermore, we show how success prediction can lead to more efficient policy termination. Standard practice in sim-to-real assembly is to execute policies for a fixed duration [7] which can lead to premature termination or delays. Instead, the policy can terminate when it believes it is in a successful state. We show that success prediction, also trained in simulation, robustly transfers to the real world and does so more reliably when using force observations [16].
@@
    A method to specify maximum allowable contact-force during policy execution. This results in policies that exhibit safe search behaviour even with significant levels of position estimation error (up to $5mm$).
@@
    A dynamics randomization scheme that reduces tuning to an interpretable scalar force-threshold parameter (instead of controller gains).
@@
    A method for success prediction that enables automatic force-threshold tuning and efficient policy termination, reducing delay times up to $66\%$.
@@
Results are shown for over $1000$ real-world trials and multiple tasks. We plan to release the code with the paper.
@@ II-A Assembly Tasks
Figure 2: FORGE is evaluated on three tasks from Factory [5]: Peg Insertion, Gear Meshing, and Nut Threading. Each task is trained solely in simulation (top) and transferred directly to the real robot (bottom).
Each task involves mating two parts: one grasped and another fixed to the workspace. For our main evaluations, we consider all three tasks from Factory [5] and demonstrate the first sim-to-real transfer for threading a small M16 nut (see Fig. 2). We also consider forceful snap-fit insertion and multi-step assembly in Sec. V-D and Sec. V-E respectively.
Peg Insertion: A round peg with $8mm$ diameter needs to be inserted into a socket with $0.5mm$ diametrical clearance.
Gear Meshing: A gear needs to be inserted onto a peg with $0.5mm$ clearance. Other gears are present and the teeth of adjacent gears must be aligned for successful meshing.
@@ II-B POMDP Formulation
We formulate our problem as a Partially Observable Markov Decision Process (POMDP) [18, 19].The goal is to learn a policy, $\pi_{\theta}(a_{t}|o_{1},\ldots,o_{t})$, that maximizes the expected return:
$J(\pi_{\theta})=\mathbb{E}_{\tau\sim p(\tau|\pi_{\theta},\Psi)}[\Sigma_{t=0}^{\infty}\gamma^{t}r_{t}]$ (1)
@@
States ($\mathcal{S}$): A state, $s_{t}\in\mathcal{S}$ consists of the pose and velocities of the end-effector (EE), fixed part, and held part: $p^{ee},p^{fixed},p^{held}\in SE(3)$ and $v^{ee},v^{held}\in\mathbb{R}^{6}$.We also include the contact force experienced by the end-effector, $F^{ee}\in\mathbb{R}^{3}$, and time-invariant information about the dynamics properties of the robot, controller, and parts (e.g., mass or joint-friction): $\Psi=(\psi_{robot},\psi_{control},\psi_{parts})$.
Observations ($\Omega$): As it is difficult to accurately estimate the full state, all our policies observe:
@@
We do not include pose or velocity of the held part because it can move in the gripper and be difficult to track.Likewise, we do not observe $\Psi$, but include the previous action, $a_{t-1}$, to help infer dynamics. See App. -A for noise models.
@@
Transition Function ($T_{\Psi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$): $T$ is parameterized by the dynamics parameters, $\Psi$ and is specified using the IsaacGym [21] simulator. The sim-to-real gap comes from the mismatch between $\Psi^{sim}$ and $\Psi^{real}$.
Observation Function ($O:\mathcal{S}\times\mathcal{A}\rightarrow\Omega$): The position of the fixed part is assumed to have up to $5mm$ error. Gaussian noise is assumed for each of the other observations.
Reward Function ($R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$): Each task is described using a keypoint reward: $R_{kp}(p^{fixed},p^{held}_{t})$ [5, 22], which is modified to account for small, threaded geometries (see App. -B for more details). We also add two discrete bonus rewards that are given when important phases of the tasks are reached: once the held part is centered on top of the fixed part and once the task is successful:
@@
We found the bonuses led to more robust learning when there is significant pose uncertainty.
@@ III FORGE: Robust Search under Uncertainty
FORGE uses on-policy RL to learn search behaviours in simulation. A force threshold (Sec. III-A) and dynamics randomization (Sec. III-B) are introduced for robust sim-to-real transfer. FORGE also introduces success prediction (Sec. III-C) for efficient termination and force threshold tuning.
@@ III-A Force Threshold
During policy execution, excessive force can cause parts to slip or become damaged.Although it may be possible to recover from small amounts of slip with the right sensors (e.g., wrist camera or tactile), we prefer to avoid these scenarios. Instead, we propose to condition the policy on a force threshold, $F_{th}$: $\pi(a|o,F_{th})$. During training, the policy is penalized if the contact force, $F^{ee}_{t}$, exceeds the threshold. Concretely, we add an additional term to the reward function:
$R_{contact\_pen}(F^{ee}_{t})=-\beta*\max(0,||F^{ee}_{t}||-F_{th}).$ (3)
Note this is related to [10] which conditions the policy on a desired force instead of an excessive force.
At deployment time, $F_{th}$ can be set or tuned based on task requirements. Most of the tasks we consider have positive clearance and do not require much force to succeed. A relatively low force threshold is sufficient to prevent slip and tuning the threshold is not necessary. For forceful insertion, nominal insertion forces are often speficied on part datasheets and the threshold should be higher than this value. We also present an automatic tuning procedure in Sec. III-C.
@@ III-B Dynamics Randomization
To successfully deploy policies trained in simulation, it is important that the trajectory distribution experienced during training is similar to what it would be when deployed: $p(\tau^{real}|\pi_{\theta},\Psi^{real})\approx p(\tau^{sim}|\pi_{\theta},\Psi^{sim})$. The difference between these distributions is usually referred to as the sim-to-real gap. This gap is usually handled by (1) system identification (Sys-ID) [23] or (2) dynamics randomization (DR) [10, 24]. The goal of Sys-ID is to tune $\Psi^{sim}$ to be close to $\Psi^{real}$. This itself is a complicated tuning procedure that may need to be redone for every new set of parts.
Instead, we follow the DR approach which learns policies that are robust to a wide range of dynamics parameters. Concretely, we optimize a version of Eq. 1 where:
@@
The integral is approximated with Monte Carlo samples from a randomization distribution. We now describe the variables that are randomized (see App. -A for values).
Controller Randomization: The controller has a large impact on contact forces. This work uses impedance-control where applied forces are computed as:
@@
First, the policy outputs a relative pose, $a_{t}$, which is applied to the fixed part’s pose to get an absolute target pose, $p^{targ}_{t}$. This pose is clipped by an action scale, $\lambda$, to ensure that the target is not too far from the EE’s current pose. As in previous work, we use critically damped gains to ensure stable controllers: $k_{d}=2\sqrt{k_{p}}$ [10, 14, 25]. The controller thus depends on two parameters which govern how much force can be commanded: $\lambda\times k_{p}$. We randomize both quantities so that the range of maximum commandable forces is in $[6.4,20.0]N$. Note that the control parameters are not included in the observations, so the policy must adjust its behavior based on force measurements. This reduces the policy’s dependence on a particular controller implementation.
Controller tuning [7] or optimization [14] is a costly and often complex procedure. Randomization [24] has the additional benefit that the policy is robust to a range of control parameters, greatly simplifying deployment.
Part Randomization: As parts slide against each other, material friction will affect lateral forces. To ensure policies can work across a range of materials, we randomize part mass and friction [24, 26, 27].
Robot Dynamics Randomization: Due to phenomena such as joint friction [24, 28], the applied force may be smaller than the commanded force. We implement a simple way to account for this: inducing a randomized dead-zone in simulation. Each episode, a dead-zone is selected for each dimension, $F^{DZ}_{i}$, where commanded forces below this value are clamped to zero: $|F^{applied}_{i}|=\max(0,|F^{targ}_{i}|-F^{DZ}_{i})$. This enables the policy to increase its target which can help apply more force when needed or reduce steady-state error.
@@
III-C Success Prediction
Although success is clearly defined in simulation where we have access to noiseless poses, it is difficult to reliably predict in the real world [16]. Consider the nut-threading task, where the distance between a successfully threaded nut and a loose nut is a fraction of a millimeter. We propose to train a success predictor which can robustly transfer from sim-to-real. Concretely, we share the weights of the policy network with the success predictor by expanding the action space of the policy to include an early termination action: $a^{ET}_{t}\in[0,1]$. To train the policy to output the correct action, we include an early termination penalty, $R^{ET}_{t}$, which penalizes incorrect success predictions: